# LMM-Assisted Breast Cancer Treatment Target Segmentation with Consistency Embedding

### v2

@@ 4.1 Dataset
To train our model, we collected internal data cohort composed of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center. For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 599 patients with their 3D CT scans and corresponding ground truth masks. To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital. The detail of dataset is described in Table 1 and Supplementary Material. We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals. This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
Table 1: Details of data for training our multi-purpose RO-LMM. “US" denotes Ultrasound. “Path" indicates Pathology.
Task Input Output Internal Dataset External Dataset
@@
Clinical Report Summarization MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
Treatment Plan Suggestion Clinical Note Treatment Plan
Plan-guided Target Segmentation Treatment Plan &\&& CT Scan Segmentation 593 79 81
4.2 Implementation details
[…]
4.3 Evaluation Metric
The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [30]. However, acknowledging potential limitations in domain specificity for the treatment plan suggestion task, we conduct expertise-based evaluations involving two clinical experts and GPT-4.0. A board-certified expert curates a total of 5 scoring rubrics for evaluation, with detailed information available in Table S9. Using the same scoring rubrics, two clinical experts assess the performance on the plan suggestion task against ground truth labels, assigning a score of 0 or 1 for each rubric separately. Inspired by the recent evaluation methodologies such as GPT-Score [13] and G-Eval [32], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning.
To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [9] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
4.4 Baseline Models
For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [14], ChatDoctor [52], Asclepius [24], and PMC-LLAMA [49]. Additionally, we include a comparison with ChatGPT [36] utilizing few-shot in-context learning for both summary and treatment plan suggestion task.
Table 2: The quantitative results for clinical note summarization using different methods. Vanilla denotes the instruction fine tuning. CI denotes confidence interval.
@@ 4.4 Baseline Models
4.5 Results
Clinical Report Summarization. We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table 2. Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals. Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning. We also perform expert analysis on the summarized reports, and the results and experimental details are deferred to Supplementary Material.
The qualitative assessment in Table 4 further highlights the effectiveness of our proposed model, providing well-organized content and consistent formatting compared to the ground truth label. While there were minor errors observed in terms of the specific distance measurement from the nipple (red), which appears to stem from the process of integrating ultrasound and MRI images, our model accurately summarizes information in a format consistent with the provided labels. In contrast, Chat-GPT exhibits issues such as omitting essential information and generating hallucinated content (red).
Treatment Plan Suggestion. As indicated in Table 2, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes. Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines. Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach. However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
Furthermore, for the treatment plan suggestion task, we comprehensively evaluate the generated plans using our proposed expertise-based rubrics as shown in Table S9. Table 3 presents the clinical expert analysis results, where R##\## represents each rubric and C##\## represents each clinical expert. For both internal and external validation, our RO-LMM variants significantly outperform the results of GPT4.0 with few-shot in-context learning due to the specific knowledge of the domain. We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals. We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$). The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
For a more detailed understanding, examples of the treatment plan suggestion task are provided in Table 4. The proposed RO-LMM-P++ (Ours) consistently yields correct answers compared to the labeled ground truth. These results demonstrate the effectiveness of our proposed CEFTune for this specific task. While Chat-GPT has shown success in providing some level of information regarding where to treat, it falls short when it comes to suggesting the specific treatment regions and dose scheme (red).
Table 3: Clinical expert analysis for treatment plan suggestion. R# represents each rubric and C# represents each clinical expert.
@@ 4.5 Results
GPT 8.8 / 6.9 / 6.2 / 5.8 / 4.2 / 31.9 8.1 / 7.3 / 6.2 / 1.5 / 8.8 / 31.9 8.5 / 8.1 / 5.0 / 1.2 / 9.2 / 31.9 8.5 / 8.1 / 6.5 / 3.8 / 9.2 / 36.2
Table 4: Example case for the clinical report summarization task and plan suggestion task compard to baseline methods. The red scripts indicate incorrect or redundant information, and inconsistent formatting compared to the label, evaluated by a board certified clinical expert.
Clinical Report Summarization Example
User
Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
##\## Input: [MRI reports ]…[Pathology reports] …
Ground Truth
[…]
Ours
[…]
Chat-GPT
[…]
Treatment Plan Suggestion Example
User
Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
##\## Input: [Clinical notes]…
Ground Truth
RT to Lt. whole breast + IMN + SCL + AXL 40.05 Gy / 15 fx $\rightarrow$ 10 Gy boost (or 48 Gy / 15 fx with IMRT).
Ours
RT to Lt. whole breast, IMN, SCL, AXL 40.05 Gy / 15 fx + 10 Gy boost (IMRT: 48 Gy / 15 fx).
Med-Alpaca
..Breast, LLO 4:00 3cm FN, IDC, icT2N1M0, stage IIB $\rightarrow$ ypT1cN1M0, stage IIA…
ChatDoctor
…The T-stage, N-stage, ER/PR/HER2/Ki67, and Lum A status of the tumor, as well as the patient’s chemotherapy and surgical history, can all be used to determine appropriate radiation …
Chat-GPT
RT to Left breast, chest wall, L axilla Conventional fractionation 40.05 Gy / 15 fx.
Table 5: Comparison of 3D target volume segmentation performance.
@@
${}^{\dagger}$modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
Table 6: Comparison of 3D target segmentation performance between unimodal and multimodal models for both overall and specific patient types.
@@
3D Target Volume Segmentation. As reported in Table 5, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting. This can be also observed in qualitative results. As shown in Figure 4(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks. Moreover, as shown in Figure 4(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows. RO-LMM-SEG without any proposed module shows reasonable performance, yet sometimes undersegments the treatment target volume, as shown in Figure 4(b). In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated.
We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type). The comparative results are presented in the Table 6 separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG. Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively. When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%. This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
@@
Figure 4: Qualitative comparison on 3D target volume segmentation task between unimodal and multimodal approaches. Red arrows indicate incorrectly segmented regions.
5 Discussion
5.1 Analysis on Strategy of Separate Expert
To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing a unified model shown in Table 8. We observe that the unified model’s performance is significantly inferior to our distinct training strategy. This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task. These results suggest that the knowledge required for each summarization and suggestion task necessitates task-aware training strategies.
@@ 5.1 Analysis on Strategy of Separate Expert
Table 7: Ablation study on adopting separate expertise for each textual task against unified strategy.
Table 8: Ablation study on CESEG with input text variation for target segmentation performance.
@@
Table 8: Ablation study on CESEG with input text variation for target segmentation performance.
5.2 Analysis on Consistency Regularization
@@
5.3 Limitation
While our study provides valuable insights on the application of radiation oncology, there are several limitations to consider. Firstly, the dataset used in this study is currently confined to patients with their initial diagnoses, primarily focusing on breast cancer. Expanding the scope of our research to include diverse patient scenarios and other cancer types is essential for enhancing the applicability of our model. Secondly, achieving detailed target volume segmentation for Simultaneous Integrated Boost (SIB) necessitates information beyond treatment plan-guided segmentation. This additional information may be obtained from alternative imaging modalities, such as MRI or PET-CT, along with ground truth masks for organ-at-risks (OARs) and gross tumor volume (GTV). Consequently, SIB is excluded from this study. Future research will focus on developing a model for SIB by incorporating a more extensive array of information.
6 Conclusion
In this work, we introduce RO-LMM, a multi-purpose foundation model tailored for radiation oncology. Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation, mirroring real-world clinical workflows. Another key contribution of this work is the introduction of consistency technique into both text and segmentation task. Results from multi-center cohort datasets confirm RO-LMM’s promising performance and noteworthy generalization capabilities across diverse tasks. These findings mark a significant stride toward developing a versatile AI model, hinting at the potential for a multi-purpose medical AI model in radiation oncology.

### v3

@@ IV-A Dataset
To train our model, we collected internal data cohort composed of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center. For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 599 patients with their 3D CT scans and corresponding ground truth masks. To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital. The detail of dataset is described in Table I and Supplementary Material. We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals. This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
IV-B Implementation details
For instruction fine-tuning for RO-LMM-S and RO-LMM-P, we use the LLaMA-2-7B-Chat [23] model as a baseline. The maximum context length is set at 4096, and the batch size is set to 2. To train RO-LMM-SEG, we employ the 3D Residual U-Net [42] using the open-source library MONAI [^2] https://monai.io/. For incorporating the textual information within the multimodel segmentation model, we use the LLaMA-2-7B-Chat model initialized with the pre-trained checkpoint of RO-LMM-P++. We set the batch size as 1. We use the AdamW [46] optimizer for all the tasks, with a learning rate of 5e-5 until reaching 3 epochs for both RO-LMM-S and RO-LMM-P, and 1e-4 until reaching 100 epochs for RO-LMM-SEG, leveraging 4 NVIDIA A6000 GPUs for each task. The hyper-parameter $\lambda$ for CEFTune is set to 1 in all of tasks, and for consistency regularization we adopt the variants of SBERT [41] which is trained on PubMed [^3] https://www.ncbi.nlm.nih.gov/pubmed/ dataset, called PubMedBERT. [^4] https://huggingface.co/NeuML/pubmedbert-base-embeddings.
TABLE I: Training data details. CRS: Clinical Report Summarization. TPS: Treatment Plan Suggestion. PTS: Plan-guided Target Segmentation. US: Ultrasound. Path: Pathology.
Task Input Output Internal External
@@ IV-B Implementation details
CRS MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
TPS Clinical Note Treatment Plan
PTS Treatment Plan &\&& CT Scan Segmentation 593 79 81
TABLE II: Quantitative comparison for clinical note summarization. Vanilla: the instruction fine tuning. CI: confidence interval.
@@
TABLE III: Clinical expert analysis for report summarization. R#: each rubric, C#:each clinical expert.
Internal Validation (N=10)
ChatGPT LLaMa2-7B RO-LMM-S
Experts R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total $r$
External Validation (N=10) $r$
IV-C Evaluation Metric and Baseline Models
The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [47]. However, acknowledging potential limitations in domain specificity for both tasks, we conduct evaluations using two clinical experts and GPT-4.0. A board-certified expert creates five scoring rubrics for plan suggestions, detailed in Table A1. Using these rubrics, the experts score each aspect of the plans, assigning 0 or 1 for each rubric separately. Similarly, for clinical report summarization, rubrics are created and the same evaluation process is followed. Inspired by the recent evaluation methodologies such as GPT-Score [48] and G-Eval [49], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning on the treatment plan task. To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [50] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [25], ChatDoctor [24], Asclepius [27], and PMC-LLAMA [26]. Additionally, we include a comparison with ChatGPT [16] for the report summarization task and GPT-4.0 for the treatment plan suggestion task using few-shot in-context learning.
TABLE IV: Clinical expert analysis for treatment plan suggestion. R#: each rubric, C#: each clinical expert.
@@ IV-C Evaluation Metric and Baseline Models
GPT-4.0 8.8 / 6.9 / 6.2 / 5.8 / 4.2 / 31.9 8.1 / 7.3 / 6.2 / 1.5 / 8.8 / 31.9 8.5 / 8.1 / 5.0 / 1.2 / 9.2 / 31.9 8.5 / 8.1 / 6.5 / 3.8 / 9.2 / 36.2
TABLE V: Comparison of 3D target volume segmentation performance.
@@
modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
TABLE VI: Comparison of 3D target segmentation performance for overall and specific patient types.
@@
IV-D Results
Clinical Report Summarization. We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table II. Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals. Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning.
Moreover, we evaluate the generated summaries using expertise-based rubrics by two clinical experts and compare them to baselines, including ChatGPT and LLaMa-2. As shown in Table III, our ROB-LMM-S model significantly outperforms all baselines in both internal and external validations, thanks to its domain-specific knowledge. Additionally, Pearson correlation ($r$) analysis reveals strong positive inter-clinician correlations ( $>0.85$ and $>0.95$ for internal and external validation, respectively), confirming the reliability of our rubrics and the clinical relevance of ROB-LMM-S. Therefore, our ROB-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
Treatment Plan Suggestion. As indicated in Table II, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes. Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines. Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach. However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
Furthermore, we comprehensively evaluate the generated plans using our expertise-based rubrics and clinical expert analysis, as detailed in Table IV. For both internal and external validation, our RO-LMM variants significantly outperform GPT-4.0 with few-shot in-context learning, as assessed by both GPT and clinical experts, due to their domain-specific knowledge. We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals. We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$). The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
@@ IV-D Results
3D Target Volume Segmentation. As reported in Table V, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting. This can be also observed in qualitative results. As shown in Figure 3(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks. Moreover, as shown in Figure 3(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows, and RO-LMM-SEG sometimes under-segments the target volume. In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated. We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type). The comparative results are presented in the Table VI separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG. Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively. When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%. This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
Inference on Public Dataset. To validate the effectiveness of our methods when AI-generated results are reused as input, we evaluated the performance by inputting each model’s report summarization results its plan suggestion task. We conducted experiments on a publicly available text-related dataset, as detailed in Table VII. Our proposed methods consistently outperformed other specialized medical LLMs and variants of GPT (such as ChatGPT and GPT-4) in both clinical report summarization and therapy plan suggestion tasks. Notably, our RO-LMM-P++ method demonstrated substantial performance enhancements, with improvements of up to 0.55 and 0.31 over other medical LLMs and GPT variants, respectively, demonstrating our model’s robustness to error accumulation during sequential generations.
TABLE VII: Quantitative comparison results for our RO-LMM’s clinical report summarization and therapy plan suggestion performance on the publicly available dataset.
Model Pubilc available dataset (N=15)
Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$ (CI) Rouge L $\uparrow$ (CI)
Clinical Report Summarization
MedAlpaca [25] 0.597 (0.502-0.683) 0.431 (0.343-0.516) 0.597 (0.502-0.683)
ChatDoctor [24] 0.644 (0.511-0.757) 0.476 (0.370-0.573) 0.644 (0.511-0.757)
Asclepius [27] 0.696 (0.658-0.733) 0.469 (0.444-0.493) 0.696 (0.658-0.733)
ChatGPT 0.765 (0.733-0.795) 0.526 (0.498-0.554) 0.765 (0.733-0.795)
Ours (RO-LMM-S) 0.795 (0.746-0.833) 0.569 (0.530-0.612) 0.795 (0.746-0.833)
Therapy Plan Suggestion
MedAlpaca [25] 0.165 (0.111-0.233) 0.085 (0.064-0.109) 0.165 (0.111-0.233)
ChatDoctor [24] 0.189 (0.147-0.233) 0.104 (0.085-0.122) 0.189 (0.147-0.233)
Asclepius [27] 0.125 (0.102-0.151) 0.074 (0.061-0.089) 0.125 (0.102-0.151)
GPT-4.0 0.390 (0.302-0.485) 0.231 (0.179-0.292) 0.390 (0.302-0.485)
Ours (RO-LMM-P++) 0.669 (0.588-0.740) 0.450 (0.374-0.525) 0.669 (0.588-0.740)
@@
Figure 3: Qualitative comparison on 3D target volume segmentation task. Red arrows indicate errors.
@@
TABLE VIII: Ablation study on adopting separate expertise for each textual task against unified strategy.
@@
TABLE IX: Ablation study on CESEG for target segmentation performance with input text variation.
@@
V Ablation Study
Analysis on Strategy of Separate Expert To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing another baseline model that unifies summarization and plan suggestion, and show the results in Table IX. We observe that the unified model’s performance is significantly inferior to our distinct training strategy. This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task. These results suggest that the knowledge required for each task necessitates task-aware training strategies.
@@ V Ablation Study
TABLE X: Inference computational complexity. RM: RO-LMM.
External Validation (N=81)
RM-S RM-P++ RM-SEG++ Analysis
Memory (GB) 13.68 13.57 17.89 Mean 14.93
Duration (sec/patient) 6.89 $\pm$ 1.68 1.75 $\pm$ 0.51 1.54 $\pm$ 0.72 Total 10.19
VI Discussion
While our study provides valuable insights into the application of radiation oncology, several limitations must be acknowledged. Firstly, the dataset is confined to patients with initial breast cancer diagnoses. Expanding to diverse patient scenarios and other cancer types is crucial for broader applicability. Secondly, radiotherapy for breast cancer in real clinical setting often requires techniques like sequential boost or simultaneous integrated boost (SIB) to deliver higher radiation doses around the tumor bed. However, accurate tumor bed delineation requires integrating information from multiple imaging modalities, which simple plan-guided segmentation alone cannot achieve. Therefore, we did not include tumor bed delineation for boost techniques in this study to simplify the problem. Future research should focus on developing a framework capable of delineating these boost volumes to aid in clinical treatment. Finally, our model did not utilize all clinical information from the actual EMR for decision-making due to token length limitations. We only used MRI, US, and pathology results, considered most critical for initial diagnosis and treatment planning. Overcoming token length constraints to integrate all EMR data will be essential for future studies.
Despite limitations, our RO-LMM demonstrates the potential to efficiently handle a broad range of clinical workflows using a single NVIDIA A6000 GPU within a 24GB memory, completing tasks in approximately 10 seconds (see Table X). This capability facilitates local implementation even with limited GPU resources and holds the promise of reducing the workload of radiation oncologists via the automation of clinical report-based plan suggestion for delineating radiation target volumes. Furthermore, our method demonstrates its effectiveness by surpassing closed-source LLMs on both a multi-center validation set and a publicly available dataset. These results suggest that our locally tuned RO-LMM-P++ holds significant potential to replace closed-source LLMs while maintaining data privacy and compliance with local data exposure regulations.
VII Conclusion
In this work, we introduce RO-LMM, a multi-purpose, comprehensive foundation model tailored for radiation oncology. Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks encompassing overall workflow of radiation oncology: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation. Another key contribution of this work is the introduction of consistency technique into both text and segmentation task. Results from multi-center cohort datasets confirm RO-LMM’s promising performance and noteworthy generalization capabilities across diverse tasks. These findings mark a significant stride toward developing a versatile AI model, hinting at the potential for a multi-purpose medical AI model in radiation oncology.